calculate the probability of the production of the aaBbccddEeFF and AaBbCcddEeFf genotypes for the following cross:

AaBbccddEeFf x AaBbCcddeeFf

MY TEACHERS ANSWER: 1/4096

I don't believe my teachers answer is correct.
Answer #1

AaBbccddEeFf x AaBbCcddeeFf

let`s divide this into 6 different crosses

Aa * Aa

A a
A AA Aa
a Aa aa

probability of getting Aa=number of Aa/total number=2/4=1/2

probability of getting aa= number of aa/total number=1/4

Bb * Bb

B b
B BB Bb
b Bb bb

probability of getting Bb= number of Bb/total number=2/4=1/2

Cc * cc

C c
c Cc cc
c Cc cc

probability of getting Cc= number of Cc/total number=2/4=1/2

probability of getting cc= number of cc/total number=2/4=1/2

dd* dd

dd

all the progenies are dd

probability of getting dd=1

Ee * ee

E e
e Ee ee
e Ee ee

probability of getting Ee= number of Ee/total number=2/4=1/2

Ff * Ff

F f
F FF Ff
f Ff ff

probability of getting Ff= number of Ff/total number=2/4=1/2

probability of getting FF= number of FF/total number=1/4

probability of getting  aaBbccddEeFF = probability of getting aa probability of getting Bb probability of getting cc probability of getting dd probability of getting Ee probability of getting FF= 1/4 1/2 1/211/21/4=1/128

probability of getting AaBbCcddEeFf = probability of getting Aa probability of getting Bb probability of getting Cc probability of getting dd probability of getting Ee ​​​​​​​ probability of getting Ff= 1/2 1/21/211/2 1/2=1/32

probability of getting aaBbccddEeFF and AaBbCcddEeFf=probability of getting  aaBbccddEeFFprobability of getting AaBbCcddEeF=1/128​​​​​​​1/32=1/4096

so the answer is 1/4096